5y^6/3v^4*9v^6/25y
It says to multiply and simplify But for the life of me I just can not understand it

Found 2 solutions by nyc_function, solver91311:
Answer by nyc_function(2741)

Use the same rules that apply to regular multiplication of fractions. You must also use the rules for division of exponents.
(5y^6/3v^4) * (9v^6/25y) =

25/5 = 5
y^6/y = y^5
9/3 = 3
v^6/v^4 = v^2
The left side fraction becomes (y^5/1) and the right side fraction becomes
(3v^2/5).
We now multiply numerator with numerator and denominator with denominator.
(y^5/1) * (3v^2/5) = (3v^2y^5)/5
Did you follow?

Answer by solver91311(24713)





Just like multiplying any other pair of fractions: Numerator times numerator and denominator times denominator:



25 is 5 times 5, so you have two factors of 5 in the denominator and one factor of 5 in the numerator. Eliminate one factor of 5 from each of the numerator and denominator.

is six factors of in the numerator and is one factor of in the denominator. Eliminate one factor of from each of the numerator and denominator.

9 is 3 times 3, so you have two factors of 3 in the numerator and one factor of 3 in the denominator. Eliminate one factor of 3 from each of the numerator and denominator.

is six factors of in the denominator and is four factors of in the numerator. Eliminate four factors of from each of the numerator and denominator.



John